### v1

# Breaking better: Imperfections increase fracture resistance in architected lattices

Abstract
Architected materials offer unique opportunities to tailor fracture properties through local structural modifications. In this study, we investigate how the failure of architected materials with triangular lattice topology is affected by the removal of individual struts, which represent well-controlled and localized imperfections. Using a combination of macroscopic mechanical testing and digital image correlation (DIC), we analyze both global response and local crack propagation. We observe that the designed imperfections do not alter the failure initiation site nor the peak tensile load but significantly increase the work to failure. DIC-based tracking reveals that this increase correlates with deviations in the crack path and may also involve mechanisms such as crack bridging or temporary pinning near defects. These results demonstrate that small, well-characterized imperfections, when properly mastered, can be harnessed to improve failure resistance and expand the design space of architected materials beyond regular, periodic structures.
Keywords: Architected materials; Additive manufacturing; Fracture characterization; Digital image correlation (DIC); Imperfections
1 Introduction
Rapid advances in additive manufacturing have enabled the fabrication of architected materials with exceptional mechanical properties, including auxetic response [1], local strengthening [2], reprogrammable stiffness [3], and enhanced toughness [4, 5]. These materials, especially periodic lattices, are attractive for applications where low density and mechanical adaptability are critical, such as in aerospace or biomedical engineering [6, 7, 8]. Among their potential advantages, fracture resistance remains one of the most desirable yet least understood features. The discrete structure of lattices introduces unique challenges and opportunities in characterizing and controlling fracture, making it essential to understand how failure initiates and propagates in these systems [9, 10, 11, 12].
Despite the limited understanding of failure mechanisms in architected materials, various topological strategies have been proposed to improve their fracture resistance while preserving the benefits of periodic structures. One approach led to the design of demi-regular tessellations with two vertex configurations to influence the global mechanical behavior [13]. However, the resulting improvement in toughness cannot outperform regular Kagome lattices with the same relative density [13]. An alternative approach introduces uniformly distributed disorder by randomizing node positions [14] or connectivity [15, 16, 17, 18], often inspired by the imperfect and hierarchical structures found in biomaterials [19, 20, 21, 22, 23, 24]. These disordered architectures deflect cracks and distribute damage effectively, but they come at the cost of eliminating long-ranged periodicity – complicating both the application of classical fracture theories and the identification of clear structure-property relationships [25].
In lattice materials, fracture emerges from the progressive failure of discrete elements, typically struts or nodes [26, 27]. While it is understood that global failure results from these local events, the connection between individual strut failure and the formation and propagation of a macroscopic crack remains difficult to establish [28]. Most models rely on strut-level failure criteria and often interpret global responses through linear elastic fracture mechanics (LEFM) [29], although the assumptions of LEFM are not always well suited for discrete architecture [30, 31, 32]. Recent numerical efforts tend to focus on the onset of failure [8], and the link between strut-scale failure and macroscopic toughness remains poorly defined, particularly in disordered lattices [16, 33]. Experimentally, this lack of understanding is even more pronounced, as measuring the sequence and location of individual failure events remains challenging. As a result, a systematic, empirical understanding of how local failure events accumulate and organize into a propagating crack in disordered lattices is still lacking.
One possible route to bridge the gap between local failure and global fracture behavior is to focus on isolated imperfections, rather than uniformly distributed disorder. Imperfections such as missing or weakened struts naturally arise from manufacturing variability [34, 35] but may also be deliberately introduced in a controlled manner. Unlike uniformly distributed disorder, which affects the structure globally [15, 16], local defects allow for targeted perturbations that preserve the periodicity of the lattice while enabling systematic investigation of their influence on crack initiation and propagation. Although previous numerical studies have characterized the toughness of both regular [36, 37] and uniformly disordered [16, 37, 38] lattices, the extent to which individual imperfections affect global fracture behavior remains unclear. A controlled approach that introduces only a few missing struts may help clarify how local structural features interact with fracture processes, and provide means to engineer the crack path and consequently the macroscopic response of lattices.
In this work, we combine macroscopic mechanical testing and optical measurements to investigate the effect of missing-strut imperfections on the fracture behavior of equilateral triangular lattice materials. We introduce local defects by removing two struts, which enables us to unambiguously link changes in mechanical response to specific imperfections. This approach allows us to distinguish the influence of these designed imperfections from that of natural manufacturing variability and to isolate their effect on crack initiation and propagation. We show that, for the tested imperfection patterns, the increase in the work to failure is systematic while the peak tensile load is preserved. By adapting “continuum” digital image correlation (DIC) to discrete lattices, we track the crack tip, extract the crack path, and identify failure events at the lattice scale. These measurements demonstrate that missing-strut imperfections cause deviations in the crack path away from the original crack plane, which increase the fracture surface and contribute to the observed increase in work to failure.
2 Fracture characterization
2.1 Specimen design with controlled imperfections
We design compact tension (CT) specimens with a triangular lattice topology, adapted from the ASTM E-1820 standard [39] (see Fig. 1a). The unit cell width $l$ is set to $4\leavevmode\nobreak\ \mathrm{mm}$ and the strut width to 0.38 mm, resulting in a relative density of $\sim$0.33. This design represents a compromise between minimizing density, maximizing printing volume, and maintaining specimen quality. In addition to regular lattice specimens, we fabricate specimens with controlled imperfections. Specifically, we introduce imperfections by removing struts from the regular lattice design, which allows for a clear distinction between controlled imperfections and naturally occurring manufacturing variations. We consider two imperfect configurations, labeled configuration A and B (Fig. 1a), each involving two missing struts but differing slightly in their (absolute and relative) locations. The low number of controlled defects enables us to assess crack sensitivity to individual defects and evaluate their effects twice within a single test.
Figure 1: Design and setup of compact tension specimens for characterizing the fracture behavior of lattices. (a) Geometry of the compact tension specimen and location of the missing struts in imperfect lattices. Configurations A and B differ in the position of one missing strut, located at varying distances from the notch root. Dimensions are given in millimeters. (b) Mechanical testing and imaging setup, which includes a LIMESS high-resolution camera for real-time imaging of the patterned specimen surfaces during fracture characterization.
[…]
2.2 Mechanical testing and imaging
The as-printed width ($w$), thickness ($b$), and pre-crack length ($a$) (see Fig. 1a) are measured for each compact tension specimen to check their agreement with the nominal dimensions, accounting for potential warping and platform removal effects. The specimens are loaded in tension using a Zwick-Roell Z100 machine at a displacement-controlled rate of 2 mm/min, with a pre-load of 4 N. The machine records the load up to a 90% peak load drop. A high-resolution LIMESS camera is placed in front of the specimen to track crack propagation over a region of interest (ROI) of approximately $85\times 60$ mm2, as shown in Fig. 1b. We lubricate the loading pins before testing to prevent stick-slip and induced rigid-body rotations of the specimen under load. The critical stress intensity factor is estimated according to the ASTM E-1820 standard [39], which provides an approximation of the lattice toughness accounting for slight variations in dimensions across samples. The mode I critical stress intensity factor ($K_{\mathrm{IC}}$) is then defined as
$K_{IC}=\frac{P_{\max}}{b\sqrt{w}}\leavevmode\nobreak\ q\left(\frac{a}{w}\right)\leavevmode\nobreak\ ,$ (1)
where $P_{\max}$ is the peak force recorded by the load cell of the testing machine, and $q$ is a function of the $a/w$ ratio specific to the specimen design [39].
3 Macroscopic properties
3.1 Failure initiation
Figure 2: Failure initiation for regular and imperfect specimens is equivalent. (a) Representative force-displacement curves for one regular specimen and one imperfect specimen of each imperfect configuration. The peak forces $P_{\mathrm{max}}$ and ultimate displacement $d_{\mathrm{u}}$ are marked by solid and open black circles, respectively. (b) Average normalized force (with 90% confidence interval) in the elastic regime at macroscopic displacements of $d_{1}=0.5$, $d_{2}=1$, and $d_{3}=1.5$ mm, as marked by vertical dashed lines in (a), based on five regular and two times four imperfect specimens (configurations A and B). (c) Critical stress intensity factor $K_{\mathrm{IC}}$, computed using Eq. (1) and normalized by $\overline{\sigma}_{\mathrm{u}}$ (the ultimate stress of base material), and $\sqrt{l}$ (with unit cell width $l$). Error bars indicate 90% confidence intervals. (d) Images of crack tip area for regular (top) and configuration B (bottom) specimens corresponding to those shown in (a), after the first struts fail. The first broken struts are highlighted with a green circle. The regular specimen shows two struts that failed simultaneously.
We examine the macroscopic response of the regular and imperfect lattices under mode I loading. Both types of lattices exhibit similar force-displacement response, as shown for a representative specimens in Fig. 2a. Initially, the response is nearly linear elastic, extending up to the maximum force, $P_{\mathrm{max}}$. Beyond this point, the force drops abruptly, which is in line with the commonly observed brittleness of lattice materials [29, 30]. However, this force drop does not immediately lead to catastrophic failure of the specimen, which withstands significant additional displacement before ultimate failure.
The elastic response of regular and imperfect lattices is quantitatively consistent, as reported in Fig. 2a. To enable a systematic comparison and account for variability due to warping, we normalize the force with the area $b(w-a)$ at various displacement levels $d$ in the quasi-linear domain. While some inconsistency in the elastic response of the lattices is observed, likely due to material variability and environmental conditions, the missing struts do not statistically affect the mechanical response up to failure initiation (Fig. 2b).
We assess the resistance to failure initiation by computing the critical stress intensity factor using Eq. (1). The average values of $K_{\mathrm{IC}}$ for both configurations of imperfect specimens are statistically equivalent and consistent with the that of regular lattices, as shown in Fig. 2c. The inspection of the images from the first failure event, associated with the force drop at $P_{\mathrm{max}}$, reveals that the same strut breaks in both regular and imperfect specimens (Fig. 2d). Specifically, the strut at the notch root is the first to fail for almost all tested specimens. Thus, we conclude that the missing struts in the configurations considered herein do not alter the crack initiation mechanism, maintaining an equivalent critical stress intensity factor.
3.2 Work to failure increase
Figure 3: Imperfect lattices exhibit significantly higher work to failure compared to regular specimens. The work to failure, corresponding to the shaded area under the force-displacement curve, is shown for the same representative examples as shown in Fig. 2a: regular specimen (blue), imperfect specimen Type A (red), and imperfect Type B (orange). (inset) Average work-to-failure for 5 regular and 8 imperfect lattices, with error bars representing 90% confidence intervals. The black triangles indicate the work-to-failure for the force-displacement curves reported in the main figure.
As previously discussed, neither the regular nor the imperfect specimens fail completely right after the peak load. To quantify the entire energy absorption capability of the lattice, we compute the work to failure, $W$, which corresponds to the integral of the force-displacement curve, given by:
$W=\int_{0}^{d_{\mathrm{u}}}P(d)\leavevmode\nobreak\ \mathrm{d}d\leavevmode\nobreak\ ,$ (2)
where $P(d)$ is the measured force at a given displacement, $d$, and $d_{\mathrm{u}}$ is the ultimate displacement at failure (see Fig. 2a). Considering representative examples of these curves for regular and imperfect specimens in Fig. 3, we observe that the imperfect lattice exhibits higher work to failure compared to the regular one.
This observation is supported by a statistical analysis of the work to failure, shown in Fig. 3 and holds for both configurations A and B. The average work to failure for imperfect lattices is approximately 30% higher than that of the regular lattices, a statistically robust finding as indicated by the 90% confidence intervals. Additionally, the confidence intervals for the imperfect specimens are notably smaller than for the regular case, despite the lower number of specimens. This reduced variability is attributed to the presence of controlled defects, which dominate failure and mitigate the influence of natural material variability inherent in the manufacturing process.
4 Local crack propagation
We attempt to correlate the observed increase in work-to-failure to changes in the local failure mechanisms by analyzing the crack path. Identifying the crack path and tip from the acquired images (Fig. 4a) is challenging due to the high slenderness of the lattice beams and the resulting variations in gray level. As an alternative, post-mortem inspection of the specimens (Fig. 4b) presents its own difficulties, as additional damage may occur during specimen removal from the testing machine. To overcome these challenges, we employ a DIC-based approach, which systematically identifies the crack morphology and tip position directly from the images acquired during mechanical testing.
4.1 DIC to unravel the crack path morphology
For DIC purposes, we pattern the specimen surface with Dupli-Color white and black spray paint. The high-resolution LIMESS camera position is adjusted to monitor a ROI that includes the notch root and the free edge of the specimen on the left side for mesh backtracking (see Appendix A). The images acquired during testing are processed using the finite-element-based DIC code Correli 3.0 [40] using a triangular mesh whose elements correspond exactly to the cells of the lattice specimen (see Fig. 4c). DIC exploits the contrast between the patterned specimen and the dark background, and minimizes the gray-level residual fields
$\rho_{\textbf{u}}(\textbf{x})=f(\textbf{x})-g(\textbf{x}+\textbf{U}(\textbf{x}))\leavevmode\nobreak\ ,$ (3)
where $f$ and $g$ are the reference and deformed images, respectively [41], and U is the in-plane displacement field (Fig. 4d), which is measured through this approach. It is worth noting that the mesh is backtracked, as detailed in Appendix A, to ensure that the elements deform during loading and the node positions remain aligned with the lattice joints (see gray lines representing the elements in Fig. 4d).
Figure 4: Automated detection of the crack path morphology by DIC for a representative imperfect specimen (configuration B). (a) Image from the LIMESS camera at the end of the mechanical test but before specimen removal illustrating a complex crack path morphology. (b) Post-mortem image of same specimen as in (a). Yellow circles indicate detected lattice damage and the orange circle marks the deduced crack tip position. (c) Region of interest (ROI) meshed with triangular elements. (d) Vertical displacements field $U_{y}$ on the deformed mesh demonstrating an accurate analysis. (e) Elementary principal strain ($\varepsilon_{1}$) field over the deformed mesh for the last acquired image shown in (a). (f) Identified crack path based on an elementary strain threshold. The centroid of the elements considered broken are marked with white dots and the crack tip is highlighted by a green marker.
To determine the crack path and tip, the displacement field obtained from DIC are post-processed. We compute the maximum principal strain ($\varepsilon_{1}$) in each element, which reveals localized regions of higher strains along the crack path at the end of the test (see Fig. 4e). To identify broken lattice cells and locate the crack tip, we apply a strain threshold of $\varepsilon_{\mathrm{th}}=0.27$ which effectively distinguishes broken from intact cells. A qualitative comparison of the original image at the end of the test (Fig. 4a), the post-mortem analysis (Fig. 4b), and the DIC-based detection (Fig. 4f) across all specimens shows good agreement, thereby indicating that the DIC-based approach reliably and with minimum effort captures the crack path morphology and tip location.
4.2 Linking crack path morphology to work to failure
Figure 5: Possible mechanisms contributing to the increased work to failure for imperfect specimens. (a) Crack path morphology for the three representative examples shown in Fig. 1. The position of each damaged lattice cell is indicated by its centroid, extracted from the DIC-based analysis. The $x$ and $y$ axes are normalized by the lattice cell width $l$ and height $h=l\sqrt{3}/2$, respectively. Lines connecting the centroids are included as visual guides only. The notch tip is marked by a gray triangle at the origin. The maximum crack path deviation $\Delta_{y}$ for configuration B is indicated by a black arrow. (b) Average work to failure (W), also reported in Fig. 3-inset, plotted as a function of the normalized maximum crack path deviation $\Delta_{y}/h$ for regular and imperfect specimens. Error bars represent 90% confidence intervals. (c-top) Image of the crack tip region for a representative configuration B specimens. (c-bottom) Corresponding schematic illustration of the crack path. Broken lattice cells are marked in blue; broken struts are shown as blue zig-zag lines, and intact struts between broken cells are highlighted in red.
To gain insight into the cause for increased work to failure in imperfect specimens, we compare the crack path morphologies obtained by DIC for the different specimen configurations. Using the three representative examples, one for each configuration, we observe that the crack path in the regular specimen remains relatively straight, with only minor deviations from a perfectly straight trajectory (see Fig. 5a). In contrast, the crack paths in the imperfect specimens show more pronounced deviations. In configuration A, the crack progressively veers away from the notch, without any major changes in direction. In configuration B, the crack initially deviates from the straight trajectory and then changes direction, resulting in a more complex path (see Fig. 5a).
To systematically analyze this effect, we introduce a simple geometric measure of crack path tortuosity, namely, the maximum crack path deviation $\Delta_{y}$, defined as the difference between the maximum and minimum $y$-positions along the crack path (Fig. 5a). By examining the correlation between the work to failure and $\Delta_{y}$, we find that configuration A specimens exhibit a consistent and statistically significant increase in crack path deviation (Fig. 5b). This observation supports the interpretation that the deviations in crack trajectory induced by the imperfections contribute to increased work to failure. In contrast, configuration B shows only a slight, statistically insignificant increase in $\Delta_{y}$, suggesting that other mechanisms may play a more prominent role in the increased work to failure for this configuration (Fig. 5b); a point that will be discussed in the following section.
5 Discussion
A key observation in this study is the increased work to failure in imperfect lattice specimens and its potential link to deviations in the crack path. In configuration A, the presence of missing struts consistently led to a more tortuous crack path, as reflected by the increased in its maximum deviation. These deviations may influence the work to failure via two mechanisms. First, they increase the effective fracture surface relative to the crack length, thereby requiring more energy to propagate the crack. Second, the modified crack trajectory may locally reduce the stress intensity at the crack tip, delaying further propagation and requiring additional work. Both effects could contribute to the observed increase in work to failure. These findings indicate that crack path modifications, induced by local imperfections, serve as a mechanism to delay or extend failure in architected materials.
While configuration A exhibited a clear correlation between crack path deviation and increased work to failure, the behavior of configuration B suggested that other mechanisms may also contribute. In these specimens, the increase in work to failure was not accompanied by a statistically robust increase in crack path deviation. One possible explanation is the presence of crack bridging [42], where some struts within the crack path remain intact for part of the failure process, temporarily sustaining load and increasing energy dissipation (Fig. 5c). Another potential mechanism is crack pinning or temporary arrest at or near the location of the defect, which could delay propagation and result in a more extended loading phase [43, 44]. These mechanisms remain hypotheses, as they are not directly confirmed by the current analysis, but they are consistent with the observed mechanical response and warrant further investigation.
In the reported experiments, the presence of a small number of missing struts did not affect the location or mechanism of crack initiation. In all tested specimens, both regular and imperfect, fracture consistently initiated at the notch root, where the stress concentration is highest. This observation indicates that, for the specific configurations considered herein, local imperfections placed away from the notch root do not significantly interfere with the initiation process. However, this result is not necessarily general. Some imperfections located closer to the notch root could modify the local stress field and potentially influence the onset of failure. Moreover, as the number of imperfections increases, the likelihood of one occurring near the notch becomes higher, which could compromise the robustness of crack initiation. These considerations highlight the need for further investigation into how the position and density of imperfections influence the early stages of fracture.
The results of this study suggest that local imperfections – typically seen as detrimental – can, in some cases, enhance fracture performance. In particular, the increased work to failure in imperfect specimens shows that small disruptions to the lattice topology can alter crack propagation while preserving strength and crack initiation. This observation points to the potential of defect engineering as a strategy to improve toughness in architected materials, echoing ideas explored in bioinspired and damage-tolerant designs where irregularities are used to guide or deflect cracks [45]. The present findings support the notion that crack path control, through deliberate confined geometric variation, may serve as a useful design principle for improving robustness in lattice-based structures.
This study focused on two specific imperfection configurations, both involving a small number of missing struts placed away from the notch root. While this choice allows for isolating and interpreting individual effects, the conclusions do not capture the full range of possible imperfection geometries or densities. In addition, the metric used to quantify crack path tortuosity is simple and does not account for the curvature or spatial complexity of the fracture trajectory. More refined geometric or energetic measures may reveal additional mechanisms or distinctions between configurations. Future work could explore a broader set of defect types, positions, and numbers to assess more systematically their influence on the fracture behavior. Combining such experimental studies with simulations could help disentangle local stress redistributions, crack pinning, and bridging effects. Extending the approach to different lattice topologies or loading modes would also clarify how general these observations are and inform design strategies for tougher, more failure-resistant architected materials.
6 Conclusion
In this work, we investigated the fracture behavior of architected materials with a triangular lattice topology by introducing controlled missing-strut imperfections and analyzing both macroscopic and local failure responses. Using a DIC-based approach, we tracked the crack path and identified the crack tip at the end of the test, providing detailed insight into local failure mechanisms. Our results showed that a small number of imperfections did not affect crack initiation or peak load but could significantly increase the work to failure. This increase was linked to deviations in the crack path, particularly in one configuration where the crack became more tortuous. In other cases, mechanisms such as crack bridging or pinning may have contributed to enhanced work to failure.
These findings suggest that local imperfections, often seen as defects, can be leveraged to improve energy dissipation during fracture. More broadly, they highlight crack path control as a potential design strategy in architected materials. Future studies should explore a wider range of imperfection types and placements, and integrate experiments with simulations to uncover the full potential of defect engineering for fracture-resistant design.
7 Acknowledgements
The authors acknowledge support from the Swiss National Science Foundation under the SNSF starting Grant (TMSGI2_211655). A.S. was supported by the Swiss National Science Foundation under grant number 200021_200343. The authors thank Dr. Mohit Pundir, Luca Michel, and Dr. Jan Van Dokkum for fruitful discussions and feedback.
8 CRediT authorship contribution statement
Alessandra Lingua: Conceptualization, Methodology, Investigation, Formal Analysis, Data Curation, Visualization, Writing – Original Draft. Antoine Sanner: Conceptualization, Writing – Review & Editing. François Hild: Methodology, Formal Analysis, Writing – Review & Editing. David S. Kammer: Conceptualization, Supervision, Writing – Review & Editing, Funding acquisition
9 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
10 Data Availability
The data are available on the ETH research collection: [DOI]
Appendices
Appendix A Mesh backtracking
Figure 4c shows a triangular mesh, referred to as the DIC mesh, which precisely aligns with the strut joint locations of the lattice. This mesh was directly generated from the STL file used for 3D-printing. To position it, a backtracking procedure was applied, which was previously introduced in Ref. [46].
The process begins with the nominal STL geometry, as shown in Fig. 1. A fine triangular mesh was then constructed based on the location of the strut joints in the STL model, and cropped to match the selected region of interest for DIC analyses (Fig. 4c). From this mesh, a mask was generated in which pixels belonging to the mesh had their gray level equal to the mean gray level of the lattice and the other pixels had their gray levels set to the mean level of the image background. This mask corresponded to the nominal configuration in which the two meshes were created. Next, a coarser auxiliary mesh enclosing the region of interest was constructed. DIC was performed between the reference image of the experiment and the mask of the nominal configuration. Once convergence was reached, the auxiliary mesh was deformed according to the estimated displacement field and expressed in the nominal frame. Knowing the nodal displacements of the auxiliary mesh, the displacement of any other point (here the centers of the strut joints) from the reference to the nominal configurations was determined via mesh interpolation. The DIC mesh was finally backtracked by applying the opposite of these interpolated displacements to its nodes, thus mapping it from the nominal to the reference configuration.